OpenStudy (anonymous):

so im doing a portfolio project and i have to write a rule for this chart that i just made which is explaining the growth of bacteria my chart is - 1hr=16 2hr=256 3hr=4096 4hr=65386 5hr=1048576 6hr=16777216 its multplies each hr. by 16. but how would i write that rule to be true. cause i dont think xtimes16 is correct

OpenStudy (solomonzelman):

you want an exponential function, really.

OpenStudy (solomonzelman):

Let's look at what exponential functions mean in general. I won't complicate by considering negative exponents, just a \(«\)real\(»\) example.

OpenStudy (anonymous):

Can you give me the rule and ill fill in the numbers. If i have the "rule" i can figure the rest out.

OpenStudy (solomonzelman):

Ok, the exponential function (the "rule" you asked for) is: \(\color{#000000}{ \displaystyle f(x)=a\cdot (b^x) }\)

OpenStudy (solomonzelman):

a - is the initial population b - the growth factor (times which you multiply every year/month/day or whatever time-frame you are using)

OpenStudy (solomonzelman):

Now, and example ...

OpenStudy (anonymous):

okay ima write it and solve it and if i think i did it right ill put it up here and you can like check it or whatever

OpenStudy (anonymous):

never mind i dont see how this works at all with the table. what would i put for the exponent

OpenStudy (solomonzelman):

EXAMPLE: Sam has two cancer cells in his body, and every day cancer cells triple. How do you represent this situation as a function of time? \(\text{----------------------------------------------------------}\) You know that before the first day has passed (i.e. when t=0), Sam has 2 cells. So, record: \(\color{#000000}{ \displaystyle t=0\quad\quad \quad y=2 }\) Next, after every new day passes, Sam's cancer cells triple. That means that you can record: \(\color{#000000}{ \displaystyle t=1\quad\quad \quad y=6 }\) Note that I said t=1, for day 1, and since 1 day has passed, I multiply the y-value times 3 (b/c the cells triple daily). Abiding this pattern that cancer triple each day, I get the following: \(\color{#000000}{ \displaystyle t=0\quad\quad \quad y=2 }\) \(\color{#000000}{ \displaystyle t=1\quad\quad \quad y=2\times 3=6 }\) \(\color{#000000}{ \displaystyle t=2\quad\quad \quad y=2\times 3^2=18 }\) \(\color{#000000}{ \displaystyle t=3\quad\quad \quad y=2\times 3^4=54 }\) and after \(t\) days, I have: \(\color{#000000}{ \displaystyle y=2\times 3^t }\) So, you can record: \(\color{#000000}{ \displaystyle y=2\times 3^t }\)

OpenStudy (solomonzelman):

it is a fairly difficult topic when first introduced, and if there are difficulties understanding, I get it. I hope tho' that this is a helpful example.